"Joan bought roses and carnations for a wedding. Roses cost $25 per dozen and carnations cost $12 per dozen. Joan spent $396 on 20 dozen flowers. How many dozen roses did she buy?"
Answer by sofiyac(983)

lets say she got x dozen roses and y dozen carnations
then
x+y=20
25x+12y=396 solve the system of equations. I"m going to use elimination method, so i'll mulitply first equation by -12 and then add it to second equation
-12x-12y=-240
25x+12y=396
-------------
13x=156 divide each side by 13
x=12
so 12 dozen roses, now we can plug that into either of the original equations and solve for y, i'm going to use the first equation
12+y=20 subtract 12 from each side
y=8 so 8 dozen carnations